# The gauge symmetry violation in the QCD ground state

### v3

@@ Abstract
The purpose of this work is to report on convincing progress made in the investigation of the true dynamical and gauge structures of the QCD ground state (vacuum). It also sheds new significant insights into a mass squared dynamical generation in it. For this purpose, we use only the tensor algebra derivation rigorous rules in dealing with the corresponding skeleton loop integrals, contributing into the full gluon self-energy. The novel constraint on any solution to QCD has been exactly and uniquely derived in the framework of the the Slavnov-Taylor identities for the gauge particles propagators. Its general solution makes it possible to prove that the gauge symmetry of the QCD Lagrangian is broken in its ground state. The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared. The exact proof has been given that it cannot be disregarded from the theory and its ground state by any means. The gauge symmetry of the QCD Lagrangian forbids such kind of terms, but the perturbative renormalizability of the theory will not be affected. The particular solution of the novel constraint describes the gluon field configurations with unbroken gauge symmetry in the QCD vacuum. The solution with broken gauge symmetry dominates over the solution with unbroken one at large distances, while being in agreement with it at short distances.
@@
Theoretical particle physics, Quantum Chromodynamics, ground state, full gluon self-energy, gauge
@@ 1 Introduction
However, this purpose remains a formidable task yet because of the multiple dynamical and topological complexities of low-energy particles (hadrons) physics, originated from QCD and its ground state (vacuum). This happens because QCD as a gauge theory of the gluon-gluon and quark-gluon strong interactions suffers from a few important conceptual problems. Here we focus on the two of them, being in the close connection with each other.
A. Whether the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is a symmetry of its ground state or not?
B. The dynamical generation of a mass squared in the vacuum of QCD, since the gauge invariance of its Lagrangian forbids such kind of terms, for example like the gluon mass term $M^{2}_{g}A_{\mu}A_{\mu}$.
The solutions of these fundamental problems is a key to the explanation of the first confinement phase transition in QCD. The properties and symmetries of the QCD Lagrangian, and thus including its Yang – Mills (YM) part, are well-known [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The propagation of gluons is one of the main dynamical effects in the QCD ground state. The importance of the corresponding equation of motion is due to the fact that its investigation is supposed to reflect the dynamical and gauge structures of the QCD ground state. The gluon Schwinger – Dyson (SD) equation [2, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] is highly non-linear (NL) because of the self-interaction of massless gluon modes. They are taken into account by the full gluon self-energy, which is the main part of the gluon SD equation itself. We will show that the general properties of the full gluon self-energy point out on some new dynamical and gauge aspects of the true QCD ground state. Due to these impacts, the novel insights into a mass dynamical generation in the QCD vacuum are also present within proposed below formalism. It is based on the tensor algebra derivation rigorous rules only, which is widely using in the theoretical and mathematical physics.
@@ 2 Gluon SD equation
The system of the SD equations of motions in QCD, including the above-mentioned gluon SD equation, in general, contains much more informations on the properties of the theory than its Lagrangian can provide at all [2]. So it might be the place where a mas squared scale parameter, forbidden by the QCD Lagrangian gauge symmetry, could be generated. Among them, the special role belongs to the equations, describing the propagation of the fundamental constituents in the QCD vacuum – gluons and quarks. The gluon SD equation has a rather complicated tensor structure because of its above-mentioned NL character. Let us emphasize that all the known interactions in nature (electroweak, gravitational and strong) are described by the gauge theories. The common mathematical language for them is the tensor algebra, which is only one used throughout this paper. That is why our works will be easily understood by a broad spectrum of researches from the different areas of the theoretical physics. The gluon SD equation analytically looks like
@@
where $g_{\rho\sigma}=T_{\rho\sigma}(q)+L_{\rho\sigma}(q)=T_{\rho\sigma}(q)+q_{\rho}q_{\sigma}/q^{2}$. In what follows the constant $\Delta^{2}_{t}(D)$ will be called as the tadpole term itself, since it deviates from the definition (2.4) by the tensor $g_{\rho\sigma}$ only.
@@
The tadpole term $\Delta^{2}_{t}(D)$, defined in eq. (2.4), is nothing else but the QD skeleton loop integral, which is already regularized from below and above as well as all the other such kind of constants which may appear in the theory by any possible ways (the subtraction schemes, the iteration solutions, etc., see below). Being a mass squared regularized quantity, it is explicitly present in the QCD ground state, see eq. (2.5). However, it violets the gauge symmetry of the QCD Lagrangian. Then a natural question arises why is it present in the vacuum of QCD at all if it makes the theory to look non-renormalizable from the very beginning? The standard solution to this dilemma was to remove this and all other QD constants in order to make theory renormalizable in accordance with the QCD Lagrangian’s exact gauge symmetry. However, here we are going to solve this dilemma by asking the question is it possible to retain the tadpole term in the QCD vacuum, but without affecting the PT renormalization properties of the theory? This is important to ask because, unlike to other QD constants, the tadpole term is only one which dynamically generates a mass squared scale parameter, dominated by the point-like four-gluon vertex only, and thus does not depending on the external gluon momentum [6]. Also, one of our primary aims will be to clarify its role in the dynamical and gauge structures of the QCD vacuum in more detail.
@@ 3 Exact constraint on any solution to QCD
It is well known that all the Slavnov-Taylor (ST) identities which appear in QCD [1] are important for the renormalizability of the theory, and thus they ”are exact constraints on any solution to QCD” [2]. Among them the ST identities [2, 24, 35] for the gauge particle propagators play a dominant role, since they determine the gauge structure of the QCD ground state (vacuum).
@@
Let us note in advance that from the very beginning the gauge-fixing parameter for the full gluon propagator $\xi$ is not to be identified with its free counterpart $\xi_{0}$. Since the gauge freedom in QCD is determined by $\xi_{0}$, there should exist a relation between them.
@@
Let us now remind that in QCD the quark skeleton loop contribution into the full gluon self-energy can be always made transverse because of the electric charge conservation which flows around the closed loop (in complete analogy with QED, see [10] and derivations below). Therefore, $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}=0$ and this relation holds independently from the gluon contribution, which has been defined in eq. (2.3). At the same time, it is well-known that in QCD that just the Faddeev – Popov (FP) ghost contribution [43], $\Pi^{gh}_{\rho\sigma}(q)$ makes the transverse relation for this term satisfied (independently from the tadpole term (2.4), which itself has the transverse projection operator structure). It is instructive to emphasize that each term in the sum (2.3) cannot be made transverse, only the whole sum. How explicitly this works can be found in any text book on QCD. For the most transparent calculations up to one loop contribution to the transverse gluon self-energy (2.3) see for example [3, 4, 5, 6]. Of course, such a transversality will be achieved in the higher loop iteration terms as well, i.,e, it is a general fact, as it is requested by the satisfied transverse relation for the (2.3) term. It is a sum of the skeleton loop contributions, as underlined from the very beginning in this work. This means that $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ and this relation holds independently from the quark contribution.
@@ 3.1 General solution
We will call it as the generalized gauge-fixing function approach to QCD as well. It has been exactly (i.e., without making any kind of the simplifications) and uniquely defined in a such new manner for the first time.This expression demonstrates the NL dependence $\xi=f(q^{2};\xi_{0})$ on $\xi_{0}$. The linear relation $\xi\rightarrow\xi_{0}$ will be recovered in the PT $q^{2}\rightarrow\infty$ limit, when the ratio $(\Delta^{2}_{t}(D)/q^{2})$ is to be suppressed in this regime at finite $\xi_{0}$. We have established the general boundary condition in QCD for the behaviour of the gauge-fixing function in the PT limit for the regularized full gluon propagator. Since the general solution (3.16) has been obtained within the ST identities formalism, we will call it as the general solution to QCD. Let us underline that to know it means to know the true gauge structure of the QCD vacuum and its theory at the fundamental quark-gluon level (for more detailed description see section 5 as well).
@@ 3.2 Particular solution
but these equalities should be put by hand in this case, i.e., they are the prescriptions rather than the exact mathematical results. We will call it as the particular solution to QCD. Let us note in advance that these conditions determine the so-called conventional or, equivalently, conventional PT QCD (though the coupling constant remains always strong apart from the asymptotic free (AF) regime [2, 38, 39]). Precisely, this solution to QCD has been studied the last five decades. In this solution the gauge symmetries of the QCD Lagrangian and its ground state coincide, like in QED (see section 6 as well).
@@ 4 Proper subtraction scheme
QCD being much more complicated quantum field gauge theory than QED, has the two satisfied transverse relations (3.14), and so requires much more careful investigation in QCD. They have to be included in a self-consistent way into the proper subtraction scheme. Finally this will lead to the removal of the corresponding QD constants from the theory. How they fixed the gauge-fixing function approach to QCD in the most general and unique ways has been just described here.
@@ 5 General solution to QCD
i.e., in this case the gauge-fixing parameter becomes the function $\xi\equiv\xi(q^{2};\xi_{0})$ and thus it is not a constant equal to $\xi_{0}$. Let us point out that the expression (5.33) coincides with the general solution (3.16), which is as it should be. So that, its asymptotic properties have been already discussed in the text after eq. (3.16). The generalized gauge-fixing function (5.33) is the known function of its arguments, while the relation (5.31) is the NL transcendental one. Behind the general inequality $\xi\neq\xi_{0}$ is the regularized constant $\Delta^{2}_{t}(D)$ as its dynamical source.
Substituting equations (5.28) and (5.29)-(5.30) into the general decomposition (3.8) for the full gluon propagator, one finally obtains
@@
This is the generalized gauge-fixing function since it depends on the constant $\Delta^{2}_{t}(D)$, and when it is zero, one recovers the gauge-fixing parameter for the free gluon propagator. In this ST identity the gauge-fixing parameter $\lambda^{-1}$ is convenient to vary continuously from zero to infinity. The functional dependence of the generalized gauge-fixing function $\xi$ (5.33) is fixed up to an arbitrary gauge-fixing parameter $\xi_{0}=\lambda^{-1}$. Unless we fix it, and thus $\xi$ itself, we will call such situation as the general gauge dependence (GGD), see equations (3.7)-(3.8), (5.32)-(5.34). Choosing $\xi_{0}=\lambda^{-1}$ explicitly, we will call such situation as the explicit gauge dependence (EGD). For example, $\xi_{0}=\lambda^{-1}=0$ is called the unitary (Landau) gauge, $\xi_{0}=\lambda^{-1}=1$ is called the t’ Hooft–Feynman gauge, etc. [30, 31, 32, 33]. The formal $\xi_{0}=\lambda^{-1}=\infty$ limit is called as the canonical gauge in [32]. This distinction seems a mere convention, but, nevertheless, it is useful one in QCD because of the presence of a mass scale parameter in its ground state. The generalized gauge requires that there is no other functional expression for $\xi$, apart from given by the relation (5.32) at finite $\xi_{0}=\lambda^{-1}$, in the full gluon propagator (5.34) and the ST identity (5.35) for the regularized gluon fields.
The system of the regularized equations (5.34)-(5.35) constitutes that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is not a symmetry of its ground state, and it has been exactly and uniquely defined. They are present in the form suitable for the NP renormalization program to be performed. The tadpole term enters the full gluon self-energy linearly, see the expression (3.12). However, in the full gluon propagator (5.34) it appears in the NL way, because its contribution has been summed up with the help of the gluon SD eq. (5.28). The term $(\Delta^{2}_{t}(D)/q^{2})$ in the expressions (5.34)-(5.34) is to be suppressed in the PT $q^{2}\rightarrow\infty$ at the finite that $\xi_{0}=\lambda^{-1}$ (the above-mentioned canonical gauge will be investigated in detail in the forthcoming paper). The full gluon propagator will behave like the free gluon propagator $\sim 1/q^{2}$. This is the one of the necessary constraints that a theory is perturbative renormalizable. Other ones such as the corresponding behaviour of the spinor Green’s function, a unitary of $S$-matrix and analyticity (causality) [2], evidently are beyond the scope of the present work. Here we have proven that the true dynamical and gauge structures of the QCD/YM ground state are much more complicated than it is required by its Lagrangian’s gauge symmetry. But despite the gauge symmetry is broken in the QCD ground state, its PT renormalizability is not affected within the QCD full gluon propagator in the generalized gauge (5.33), i.e., it has the PT renormalizable behaviour at large $q^{2}$.
@@
It is necessary to underline that none of the prescriptions (i.e., putting some of them zero by hand) have been introduced and none of the truncations, approximations and assumptions have been done as well in the obtaining of these relations. In other words, all of them are exact mathematical results in the general solution to QCD.
6 Particular solution to QCD
The widely used up to present days the linear relation $\xi=\xi_{0}$ is only a particular solution of the derived here the exact constraint (3.15). As mentioned above in section 3, it is defined when the tadpole term has to be put zero by hand, i.e., $\Delta^{2}_{t}(D)=0$, then $\xi=\xi_{0}=\lambda^{-1}$ as well and vice versa. The corresponding expressions for the full gluon propagator and the ST identity in the particular solution to QCD can be independently derived absolutely in the same way as has been derived the corresponding expressions for the general solution to QCD in the previous section. However, the simplest way to show the particular solution is to put formally $\Delta^{2}_{t}(D)=0$ in the system of eqs. (5.34)-(5.35), then it looks like

### v4

# Confining QCD theory and Mass Gap

@@ Abstract
The novel constraint on any solution to QCD has been exactly and uniquely derived in the framework of the the Slavnov-Taylor identities for the gauge particles propagators. It has the two independent solutions. 1). The first one leads to the general QCD, possessing the confinement phase transition. 2).The second one leads to the particular QCD, or, equivalently, the conventional QCD, does not possessing it. The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared, contributing to the full gluon self-energy. Its renormalized version is conventionally called a mass gap. By formulating the proper subtraction scheme, we prove that it cannot be disregarded from the theory and its ground state by any means, but the perturbative renormalizability of QCD will not be affected.The general QCD forbids the free gluons to appear in the physical spectrum at large distances, and, at the time, has an asymptotic freedom behaviour at high energies. Finally, we describe how our advance results are connected to the Jaffe-Witten’s theorem.
@@
Theoretical particle physics, Quantum Chromodynamics, vacuum, full gluon self-energy, gauge, mass gap
@@ 1 Introduction
However, this purpose remains a formidable task yet because of the multiple dynamical and topological complexities of low-energy particles (hadrons) physics, originated from the QCD and its ground state (vacuum). This happens because QCD as a gauge theory of the gluon-gluon and quark-gluon strong interactions suffers from a few important conceptual problems. The conventional QCD, which has been investigated the last five decades, cannot explain the colour confinement problem, i.e., why the coloured objects– gluons and quarks– cannot appear at low-energies/large distances, which is an experimental fact. Secondly, it cannot explain the asymptotic freedom (AF) phenomenon, i.e., a scale breaking at high-energies/short distances [38, 39], which is also an experimental fact. To explain the both experimental facts at low- and high-energies it is necessary to answer a long-standing question how does a mass come out of a massless theory (the mass gap) [2]? even at the fundamental quark gluon level. For example, the linear rasing potential between heavy quarks requires the mass scale parameter – the string tension. Also, a scale breaking in AF regime needs a mass squared scale parameter. So that, the goal arises how to solve the above-mentioned problems, and thus to formulate the confining QCD theory, and, at same time, explaining AF phenomenon as well. It may come from the investigation of the true gauge and dynamical structures of the QCD vacuum in more detail.
The properties and symmetries of the QCD Lagrangian, and thus including its Yang – Mills (YM) part, are well-known [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The propagation of gluons is one of the main dynamical effects in the QCD ground state. The importance of the corresponding equation of motion is due to the fact that its investigation is supposed to reflect the dynamical and gauge structures of the QCD ground state. The gluon Schwinger – Dyson (SD) equation [2, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] is highly non-linear (NL) because of the self-interaction of massless gluon modes. They are taken into account by the full gluon self-energy, which is the main part of the gluon SD equation itself. We will show that the general properties of the full gluon self-energy point out on some new dynamical and gauge aspects of the true QCD ground state. Due to these impacts, the novel insights into a mass dynamical generation in the QCD vacuum are also present within proposed below formalism. It is based on the tensor algebra derivation rigorous rules only, which is widely using in the theoretical and mathematical physics.
@@ 2 Gluon SD equation
The system of the SD equations of motions and the corresponding Slavnov-Taylor (ST) identities in QCD, in general, contains much more informations on the properties of the theory than its Lagrangian can provide at all [2]. So it might be the place where a mas squared scale parameter could be generated. Among them, the special role belongs to the equations, describing the propagation of the fundamental constituents in the QCD vacuum – gluons and quarks. The above-mentioned gluon SD equation has a rather complicated tensor structure because of its NL character. Let us emphasize that all the known interactions in nature (electroweak, gravitational and strong) are described by the gauge theories. The common mathematical language for them is the tensor algebra, which is only one used throughout this paper.That is why our work will be easily understood by a broad spectrum of researches from the different areas of the theoretical physics. The gluon SD equation analytically looks like
@@
where $g_{\rho\sigma}=T_{\rho\sigma}(q)+L_{\rho\sigma}(q)=T_{\rho\sigma}(q)+q_{\rho}q_{\sigma}/q^{2}$. In what follows the constant $\Delta^{2}_{t}(D)$ will be called as the tadpole term itself, since it deviates from the definition (2.4) by the tensor $g_{\rho\sigma}$ only. So that, the tadpole term equally contributes to the transverse and longitudinal components of the full gluon self-energy (2.2).
@@
The constant tadpole term $\Delta^{2}_{t}(D)$, defined in eq. (2.4), is nothing else but the QD, i.e, UV divergent skeleton loop integral, which is already regularized from below and above, as well as all the other such kind of constants which may appear in the theory by any possible ways (the subtraction schemes, the iteration solutions, etc., see below). Being a mass squared regularized quantity, it is explicitly present in the QCD ground state, see eq. (2.5). Then a natural question arises why is it present in the vacuum of QCD at all if it makes the theory to look non-renormalizable from the very beginning? Let us remind that the theory possessing the QD quantities is not the PT renormalizable one since the logarithmic-type divergences only can be accounted for the PT. The standard solution was to remove this and all other QD constants in any possible gauge-invariant way from the theory in order to make it renormalizable in the PT sense. However, here we are going to ask the question is it possible to retain the tadpole term in the QCD vacuum, but without affecting the PT renormalizability of the theory? This is important to ask because, unlike to other QD constants, the tadpole term is only one which dynamically generates a mass squared scale parameter, dominated by the point-like four-gluon vertex only, and thus does not depending on the external gluon momentum at all [6]. Also, we will clarify its role in the dynamical and gauge structures of the QCD ground state in more detail.
@@ 3 Exact constraint on any solution to QCD
It is well known that all the ST identities which appear in QCD are important for the renormalizability of the theory, and thus they ”are exact constraints on any solution to QCD” [2]. Among them the ST identities [2, 24, 35] for the gauge particle propagators play a dominant role, since they determine the gauge structure of the QCD vacuum.
@@
It is important to note in advance that from the very beginning the gauge-fixing parameter for the full gluon propagator $\xi$ is not to be identified with its free counterpart $\xi_{0}$. Since the gauge freedom in QCD is determined by $\xi_{0}$, there should exist a relation between them.
@@
Let us now remind that in QCD the quark skeleton loop contribution into the full gluon self-energy can be always made transverse because of the electric charge conservation which flows around the closed loop (in complete analogy with QED, see [10] and derivations below). Therefore, $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}=0$ and this relation holds independently from the gluon contribution, which has been defined in eq. (2.3). At the same time, it is well-known that in QCD just the Faddeev – Popov (FP) ghost contribution [42] $\Pi^{gh}_{\rho\sigma}(q)$ makes the transverse relation for this term satisfied (independently from the tadpole term (2.4), which itself has the transverse projection operator structure). It is instructive to emphasize that each term in the sum (2.3) cannot be made transverse, only the whole sum. How explicitly this works can be found in any text book on QCD. For the most transparent calculations up to one loop contribution to the transverse gluon self-energy (2.3) see for example [3, 4, 5, 6]. Of course, such a transversality will be achieved in the higher loop iteration terms as well, i.,e, it is a general fact, as it is requested by the satisfied transverse relation for the (2.3) term. It is a sum of the skeleton loop contributions, as underlined from the very beginning in this work. This means that $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ and this relation holds independently from the quark contribution.
@@ 3.1 General solution
This relation has been exactly (i.e., without making any kind of the simplifications) and uniquely defined in a such new manner for the first time.This expression demonstrates the NL dependence $\xi=f(q^{2};\xi_{0})$ on $\xi_{0}$. The linear relation $\xi\rightarrow\xi_{0}$ will be recovered in the PT $q^{2}\rightarrow\infty$ limit, when the ratio $(\Delta^{2}_{t}(D)/q^{2})$ is to be suppressed in this regime at finite $\xi_{0}$. We have established the general boundary condition in QCD for the behaviour of this relation in the PT limit for the regularized full gluon propagator. Since the exact solution (3.16) has been obtained within the ST identities general formalism, we will call it as the general solution to QCD and its vacuum, i.e., at the fundamental quark-gluon level (see section 5 for detailed description).
@@ 3.2 Particular solution
but these equalities should be put by hand in this case. They are the prescriptions rather than the exact mathematical results. In what follows we will call it as conventional or, equivalently, the PT QCD as well (though the coupling constant remains always strong apart from the asymptotic free (AF) regime [2, 38, 39]). For its description in some more details see section 6.
@@ 4 Proper subtraction scheme
QCD being much more complicated quantum field gauge theory than QED, has the two independent satisfied transverse relations (3.14), and so requires much more careful investigation in QCD. They have to be included in a self-consistent way into the proper subtraction scheme. Finally this will lead to the removal of the corresponding QD constants from the theory. How they fixed the relations (3.16) and (3.17) in the most general and unique ways has been just described above.
@@ 5 General QCD
i.e., in this case the gauge-fixing parameter becomes the function $\xi\equiv\xi(q^{2};\xi_{0})$ and thus it is not a constant equal to $\xi_{0}$. Let us point out that the expression (5.33) coincides with the general solution (3.16), which is as it should be. So that, its asymptotic properties have been already discussed in the text after eq. (3.16). The function (5.33) is the known function of its arguments, while the relation (5.31) is the NL transcendental one. Behind the general inequality $\xi\neq\xi_{0}$ is the regularized constant $\Delta^{2}_{t}(D)$ as its dynamical source.
Substituting equations (5.31) and (5.33) into the general decomposition (3.8) for the full gluon propagator, one finally obtains
@@
The ST identity (5.35) depends on the constant $\Delta^{2}_{t}(D)$, and when it is zero, one recovers the gauge-fixing parameter for the free gluon propagator. In this ST identity the gauge-fixing parameter $\lambda^{-1}$ for the free gluon propagator is convenient to vary continuously from zero to infinity. The functional dependence of $\xi$ (5.33) is fixed up to an arbitrary gauge-fixing parameter $\xi_{0}=\lambda^{-1}$. Unless we fix it, and thus $\xi$ itself, we will call such situation as the generalized gauge dependence (GGD), see equations (3.7)-(3.8), (5.32)-(5.35). Choosing $\xi_{0}=\lambda^{-1}$ explicitly, we will call such situation as the explicit gauge dependence (EGD). For example, $\xi_{0}=\lambda^{-1}=0$ is called the unitary (Landau) gauge, $\xi_{0}=\lambda^{-1}=1$ is called the t’ Hooft–Feynman gauge, etc. [30, 31, 32, 33]. The formal $\xi_{0}=\lambda^{-1}=\infty$ limit is called as the canonical gauge in [32]. This distinction seems a mere convention, but, nevertheless, it is useful one in QCD because of the presence of a mass scale parameter in its ground state. The generalized gauge requires that there is no other functional expression for $\xi$, apart from given by the relation (5.33) at finite $\xi_{0}=\lambda^{-1}$, in the full gluon propagator (5.34) and the ST identity (5.35) for the regularized gluon fields. For the renormalized version of the ST identity (5.35) see section 6 below.
The system of the regularized equations (5.34)-(5.35), explicitly depending on the tadpole term, are present in the form suitable for the NP renormalization program to be performed. The tadpole term enters the full gluon self-energy linearly, see the expression (3.12). However, in the full gluon propagator (5.34) it appears in the NL way, because its contribution has been summed up with the help of the gluon SD eq. (5.28). The term $(\Delta^{2}_{t}(D)/q^{2})$ in the expressions (5.34)-(5.35) is to be suppressed in the PT $q^{2}\rightarrow\infty$ limit at the finite $\xi_{0}=\lambda^{-1}$ (the above-mentioned canonical gauge will be investigated in detail in the forthcoming paper). Then the full gluon propagator will behave like the free gluon propagator $\sim 1/q^{2}$ in this limit. This is the one of the necessary constraints that a theory is perturbative renormalizable. Other ones such as the corresponding behaviour of the spinor Green’s function, a unitary of $S$-matrix and analyticity (causality) [2], evidently are beyond the scope of the present work. Here one can conclude that the true dynamical and gauge structures of the QCD/YM ground state are much more complicated than it follows from its Lagrangian’s formulation. However, its PT renormalizability is not affected within the QCD full gluon propagator in the generalized gauge (5.34), i.e., it has the PT renormalizable behaviour at large $q^{2}$.
@@
It is necessary to underline that none of the prescriptions (i.e., putting some of them zero by hand) have been introduced and none of the truncations, approximations and assumptions have been done, as well as no gauge fixing by hand, in the obtaining of these relations. In other words, all of them are exact mathematical results in the general solution to QCD.
6 Renormalized ST identity and mass gap
Since the novel constraint derived in section 3 has been mainly based on the ST identities formalism, it is possible to perform the NP renormalization program for the ST identity (5.35) itself. It is instructive to present (5.35) in terms of $\xi$ and $\xi_{0}$ again, then we have
Let us now introduce the NP renormalization constant for the tadpole term as follows: $\Delta^{2}_{t}(D)=Z_{\Delta}\Delta^{2}$, where $\Delta^{2}$ is finite and positive, by definition, while the NP renormalization constant $Z_{\Delta}$ depends on all the unphysical parameters, such as regulating parameters, mentioned above, etc. In principle, the string tension in the linear rising potential between heavy quarks has to be somehow related to the finite $\Delta^{2}$, since it is connected to the transverse projection operator structure, as underlined above. It is well-known that the confining ansatz $D_{\mu\nu}(q)\sim T_{\mu\nu}(q)\Delta^{2}/(q^{2})^{2}$ just leads to a such potential. So that, the ST identity for the renormalized full gluon propagator becomes
where the renormalized gauge-fixing parameters are defined as follows: $\tilde{\xi}_{0}=Z_{\Delta}\xi_{0},\ \xi^{R}=Z_{\Delta}\xi$ and $D^{R}_{\mu\nu}(q)=Z_{\Delta}D_{\mu\nu}(q)$, i.e., the ST identity (6.39) is expressed in terms of the finite quantities only. The correct PT limit is maintained, of course. The renormalized version of the tadpole term $\Delta^{2}$ can be considered as a scale determining the NP structure of the general QCD. Conventionally, we call it as a mass gap and see discussion in section 10 as well. Such a finite mass gap $\Delta^{2}$ will appear in the full gluon propagator as well, i.e., not only in its longitudinal part (see sections 8 and 9).
Let us now demonstrate the interesting feature of the generalized gauge (6.39). It provides the self-consistency condition for the gauge choice in QCD as follows:
[…]
which is only satisfied at $q^{2}=0$, i.e., there is no any condition for the gauge-fixing parameter. The formulated self-consistency condition (6.40) points out on the inconsistency of the canonical gauge $\tilde{\xi}_{0}=\infty$ in QCD. Briefly, if everything is expressed in the terms of the finite quantities then the formal canonical gauge is forbidden to use in QCD, according to our approach to this theory.
7 Conventional QCD
The widely used up to present days the linear relation $\xi=\xi_{0}$ is only a particular solution to the exact constraint (3.15). As mentioned above in section 3, it is defined when the tadpole term has to be put zero by hand, i.e., $\Delta^{2}_{t}(D)=0$, then $\xi=\xi_{0}=\lambda^{-1}$ as well and vice versa. The corresponding expressions for the full gluon propagator and the ST identity in this solution to QCD can be independently derived absolutely in the same way as has been derived the corresponding expressions for the general solution to QCD in the previous section. However, the simplest way to show the conventional solution is to put formally $\Delta^{2}_{t}(D)=0$ in the system of eqs. (5.34)-(5.35), then it looks like